Grade 11General Physics

A paticle is executing SHM with time period T . Starting from mean position, time taken by it to complete 5/8 oscilattions is

Kaustubh Dighe
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
(5/8)T = T/4  + T/4 + T/8
for T/8 the amplitude would be A/2
Therefore,
A/2 = Asin(wt)
t = T/12
total time => T/12 + T/4 + T/4 => (T + 3T + 3T)/12 => 7T/12